# Realistic quantum network simulation for experimental BBM92 key distribution

### v1

@@ VI BBM92 simulation and theory via repeaters
For entanglement generation protocols with deterministic swapping, there exists an exact analytical expression for the expected time to generate entanglement between Alice and Bob. The average time required to generate entanglement between the end nodes of a repeater network with $2^{n}$ elementary links and with an elementary link distance of $L_{0}$ is PhysRevA.83.012323
@@
where $c$ is the speed of light and $P_{0}$ is the probability of establishing entanglement across an elementary link. $P_{0}=\frac{1}{2}\eta_{\text{link}}$, where $\eta_{\text{link}}$ is the transmission efficiency from one node to the Bell state measurement station (BSM) in the center of the elementary link, to the other node. For simplicity, we assume the photons travel in a vacuum, but to consider another material, the speed of light can be divided by the index of refraction of the material. Note the distance between a node and a BSM is $L_{0}/2$, and the total end-to-end length is $L=L_{0}2^{n}$. There are two end nodes and $r=L/L_{0}-1$ repeater nodes. Additionally, for the elementary link entanglement generation, the Bell state measurement efficiency is one half, the efficiency of a standard linear optical Bell state measurement. The minimum classical communication time to determine the entanglement generation attempt success after each attempt is also incorporated.
Figure 6: Simulated (1000 shots) secure key rates for a quantum repeater network with different numbers of repeater nodes, for different end-to-end losses, and for different elementary link Werner state fidelities $F_{i}$. The dotted lines show theoretical calculations while points with error bars show AQNSim simulations. The red lines are repeater networks with 20 dB end-to-end loss, while blue lines are repeater networks with 5 dB end-to-end loss, with darker to lighter lines representing increasing numbers of repeater nodes [0, 1, 2, 3].
@@
Applying our simulator to longer repeater networks and to the specific use case of BBM92, we simulate entanglement generation rates and predict secret key rates for a repeater network given deterministic swapping but imperfect elementary entanglement link fidelities (see Figure 6). The theory curves use the expression for the final fidelity $F$ of an end-to-end entangled pair in a repeater network generated from elementary links with entangled states which are Werner states with fidelity $F_{i}$ Briegel1998
@@
In Figure 6, we analyze entanglement generation rates given deterministic swapping in order to validate our simulations against theory. While there is no exact expression for the expected entanglement generation rate for repeater networks with probabilistic swapping PhysRevA.100.032322 , this as well as purification protocols can also be efficiently simulated with AQNSim.
Figure 6 provides insight into tradeoffs inherent in repeater networks. As the elementary link fidelity degrades, the secure key rate decreases, and at different rates for different repeater node numbers $r$. The optimum number of repeater nodes $r$ should be chosen based on the total end-to-end loss and the elementary link Werner state fidelity.
Although the simulations plotted in Figure 6 are kept general and omit implementation-specific details including detector jitter and probabilistic swapping, it is straightforward to scale up the model in complexity and detail because of AQNSim’s high degree of configurability. AQNSim can also simulate quantum gate errors and delays, user-specified classical signal delays, other noise models beyond depolarizing noise, finite quantum memory times, entanglement purification protocols, and other realistic details and extensions.
@@ VII Conclusions
We have employed a discrete event simulation framework for quantum networks implemented using the python package AQNSim, and shown it can accurately simulate key rates and QBERs for BBM92 key exchange. We have experimentally conducted BBM92 key exchange, reproduced the experimental setup with AQNSim, and shown our simulation results and experimental results align. Additionally, we have used AQNSim to solve for quantum network metrics in more complex repeater scenarios where experimental implementations do not yet exist, verifying our simulations against simple but exact theoretical models. In future work, AQNSim can be employed to simulate repeater protocols with increasing levels of realism, including protocol-specific classical message delays, quantum memories with finite storage times, chromatic dispersion and detector jitter, detector dark counts, and probabilistic swapping.
Accurate, scalable, and modular quantum network simulators are critical tools for designing quantum networks and predicting protocol performance – especially as network complexity scales to the point where theoretical expressions are lacking or are otherwise impractical to derive. The customizable level of abstraction for quantum network components in AQNSim makes it a highly versatile simulation software. As quantum technology continues to mature, our simulator and other discrete event quantum network simulators will accelerate real-world quantum secure communication deployments and advance related quantum applications including quantum sensing and distributed quantum computing.
@@
All experimental data and code other than the source code for the proprietary AQNSim python package are available at https://github.com/aliro-technologies/realistic-quantum-network-simulation.
@@ Appendix A Equipment and synchronization
Synchronization between detector pairs is necessary before timestamps and bases can be exchanged in a BBM92 protocol. Synchronization is performed during setup calibration by finding the displacement of $g^{(2)}$ peaks for different detector pairings between Alice and Bob, and solving for the delays.
@@ Appendix B Measurement correlation matrix
Figure 7: Correlation matrix calculated from experimental BBM92 data displaying the probabilities of Alice and Bob measuring a given outcome in a given basis. The coincidence window is 2 ns and other relevant parameters are provided in Table 1.
The security of BBM92 is dependent upon the orthogonality of the two mutually unbiased measurement bases that Alice and Bob use to perform measurements. We quantify the bases orthogonality in our BBM92 experiment by plotting the probabilities of correlation for each measurement outcome and basis for Alice and Bob (see Figure 7). The measurement correlation matrix for the Bell state |$\ket{\Psi^{+}}$ has perfect correlations in the XX and ZZ subspaces, and no correlations in the XZ and ZX subspaces freqbinqkd .
@@ Appendix C Parameter estimation from experimental data
The probability of error due to optical imperfections, $p_{0}$, is estimated as the experimentally calculated QBER at coincidence windows of 1 ns, where the ratio of accidental coincidences to true pair coincidences is negligible.

### v3

@@ VI Comparing Simulation and Theory for BBM92 in Quantum Repeaters
For entanglement generation protocols with deterministic swapping, there exists an exact analytical expression for the expected time to generate entanglement between Alice and Bob. The average time required to generate entanglement between the end nodes of a repeater network with $2^{n}$ elementary links and with an elementary link distance of $L_{0}$ is [7]
@@
where $c$ is the speed of light and $P_{0}$ is the probability of establishing entanglement across an elementary link. $P_{0}=\frac{1}{2}\eta_{\text{link}}$, where $\eta_{\text{link}}$ is the transmission efficiency from one node to the Bell state measurement station (BSM) in the center of the elementary link, to the other node. For simplicity, we assume the photons travel in a vacuum, but to consider another material, the speed of light can be divided by the index of refraction of the material. Note the distance between a node and a BSM is $L_{0}/2$, and the total end-to-end length is $L=L_{0}2^{n}$. There are two end nodes and $r=L/L_{0}-1$ repeater nodes. Additionally, for the elementary link entanglement generation, the Bell state measurement efficiency is one half, the efficiency of a standard linear optical Bell state measurement. The minimum classical communication time to determine whether the entanglement generation attempt was successful after each attempt is also incorporated.
Figure 6: Simulated (1000 shots) secure key rates for a quantum repeater network with different numbers of repeater nodes, for different end-to-end losses, and for different elementary link Werner state fidelities $F_{i}$. The dotted lines show theoretical calculations while the error bars show AQNSim simulations. The red lines are repeater networks with 20 dB end-to-end loss, while blue lines are repeater networks with 5 dB end-to-end loss, with darker to lighter lines representing increasing numbers of repeater nodes [0, 1, 2, 3].
@@
Applying our simulator to longer repeater networks and to the specific use case of BBM92, we simulate entanglement generation rates and predict secret key rates for a repeater network given deterministic swapping but imperfect elementary entanglement link fidelities (see Figure 6). The theory curves use the expression for the final fidelity $F$ of an end-to-end entangled pair in a repeater network generated from elementary links with entangled states which are Werner states with fidelity $F_{i}$ [9]
@@
Figure 6 provides insight into tradeoffs inherent in repeater networks. As the elementary link fidelity degrades, the secure key rate decreases, and it decreases at different rates for different repeater node numbers $r$. The optimum number of repeater nodes $r$ should be chosen based on the total end-to-end loss and the elementary link Werner state fidelity.
Although the simulations plotted in Figure 6 are kept general and omit implementation-specific details including detector jitter, it is straightforward to scale up the model in complexity and detail. In future work, one can incorporate quantum gate errors and delays, user-specified classical signal delays, other noise models beyond depolarizing noise including time-varying noise models, finite quantum memory times, entanglement purification protocols, and other realistic details and extensions which may not be readily addressed purely using analytical theory.
There is no exact expression for entanglement generation rates in repeater networks given probabilistic swapping [31]. Future work can also extend the deterministic swapping entanglement generation simulations discussed in this section to address entanglement generation in repeater networks given probabilistic swapping.
@@ VII Conclusions
We have employed a discrete event simulation framework for quantum networks implemented using the Python package AQNSim, and shown it can accurately simulate key rates and QBERs for BBM92 key exchange. We have experimentally conducted BBM92 key exchange, reproduced the experimental setup with AQNSim, and shown our simulation results and experimental results align with a smaller mean squared error than experiment and analytical theory. Additionally, we have used AQNSim to solve for quantum network metrics in more complex repeater scenarios where experimental implementations do not yet exist, verifying our simulations against simple but exact theoretical models.
We use AQNSim as a reference implementation of a discrete event simulation framework. Comparable quantum network Python simulation libraries exist and may be used for alternative implementations of the simulations presented in our work, provided they support the required features (see Appendix D).
In future work, discrete event simulation software with the requisite features and capabilities can be employed to simulate repeater protocols with additional complexity beyond that which is addressed in this work and with increasing levels of realism, including protocol-specific classical message delays, quantum memories with finite storage times, chromatic dispersion and detector jitter, detector dark counts, and probabilistic swapping.
Accurate, scalable, and modular quantum network simulators are critical tools for designing quantum networks and predicting protocol performance – especially as network complexity scales to the point where theoretical expressions are lacking or are otherwise impractical to derive. As quantum technology continues to mature, discrete event quantum network simulators will accelerate real-world quantum secure communication deployments and advance related quantum applications including quantum sensing and distributed quantum computing.
@@
All experimental data and code other than the source code for the proprietary AQNSim Python package are openly available [27].
@@ Appendix A Equipment and synchronization
Temporal synchronization between detector pairs is necessary before timestamps and bases can be exchanged in a BBM92 protocol. Temporal synchronization is performed during setup calibration by finding the displacement of $g^{(2)}$ peaks for different detector pairings between Alice and Bob, and solving for the delays.
@@ Appendix B Measurement correlation matrix
Figure 7: Correlation matrix calculated from experimental BBM92 data displaying the probabilities of Alice and Bob measuring a given outcome in a given basis. The coincidence window is 2 nanoseconds and other relevant experimental parameters are provided in the main text.
[…]
@@ Appendix C Parameter estimation from experimental data
Table 2: Comparison of different discrete event quantum network Python libraries by the features necessary for the simulations presented in the main text of this work. ✓indicates support for feature as found in the simulator’s source code or documentation; X indicates no support for feature.
AQNSim SeQUeNCe QuISP NetSquid
Continuous-wave polarization-
entangled photon source
X X
Beam splitter component X X
Waveplate component X X X
Link photon loss
Detector dead time
Detection timing jitter X
Detector dark counts
Parallelization option
for many simulation runs
X
Figure 9: (a) Entanglement creation between two end nodes (Alice and Bob), with no repeater nodes ($r=0)$, and a single Bell state measurement station (BSM). (b) The average entanglement generation rate (in units of $c/L_{0}$) for different elementary link loss values calculated from simulation (using SeQUenCE with 1000 shots) and from theory.
The probability of error due to optical imperfections, $p_{0}$, is estimated as the experimentally calculated QBER using a coincidence window of 1 nanosecond, such that the ratio of accidental coincidences to true pair coincidences is negligible. This QBER is effectively a measurement of the entangled photon source and optical setup visibility, which is a visibility that incorporates quantum state errors due to any defects in the measurement optics and integrated source and due to any imperfections in the polarization drift compensation. This estimate of the source visibility is corroborated by independently taken visibility measurements of the entire optical setup, including the entangled photon source and the measurement optics.
Appendix D Additional quantum network simulation platforms
We use the discrete event quantum network simulation Python library AQNSim for the simulations presented in our work. There are many other existing discrete event quantum network simulation Python libraries, many with comparable features and functionalities [25-28]. The simulation framework presented in this paper may, in principle, be implemented with another discrete event quantum network simulator with the necessary features. Table 2 lists the features needed to reproduce all simulations in this work while noting their presence in quantum network simulators AQNSim, SeQUeNCe [25], QuISP [26], and NetSquid [27]. As the table shows, many of the features necessary for our work are common across the other listed simulators. However, exactly replicating all of the simulations presented in this work would require custom implementation of new components or features if using a discrete event quantum network simulator other than AQNSim.
SeQUeNCe is a commonly used quantum network discrete event simulator which is open-source with several basic built-in protocols and a sizable component library. To demonstrate another example of validation for another quantum network simulation platform, we constructed a basic entanglement distribution simulation using SeQUeNCe. For the simulation, Alice and Bob are each located at a distance $L_{0}/2$ from a central Bell state measurement station and each generate shared entanglement across their quantum memories using a simple entanglement generation protocol (see Figure 9a). The average entanglement generation rates are plotted in Figure 9b against the total link loss (total loss from Alice to the BSM to Bob). We additionally incorporate the minimum classical communication time to determine entanglement generation attempt success in the entanglement generation rates. We see agreement of simulation with theory using the same analytical model used for the simulations shown in Figure 5b in our main text, here with repeater node number $r=0$.
We note that the SeQUeNCe implementation for a Bell state measurement between two photonic qubits entangled with quantum memories, SingleAtomBSM, as of the writing of this paper, applies the channel loss only for one beam splitter input arm, either Alice to BSM or Bob to BSM, when only a single photon is measured. In this work, we model entanglement generation protocols in which a photon is transmitted along each channel towards the beam splitter and each photon independently experiences channel loss. To this end, we implemented a modified version of the SingleAtomBSM class to apply channel loss independently for both arms, Alice to BSM and Bob to BSM.